OpenStudy (anonymous):

Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form.

OpenStudy (dumbcow):

first find slope \[m = \frac{y2-y1}{x2-x1}\] then you point-slope form \[y-y1 = m(x-x1)\]

OpenStudy (anonymous):

Thanks

OpenStudy (dumbcow):

your welcome i can check your answer for you too if you want

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

i got 7/11 for finding the slope , am I right ?

OpenStudy (dumbcow):

no you should get 12/6 = 2 remember with any point x is 1st number, y is 2nd number

OpenStudy (dumbcow):

(3,4) --> x1 = 3, y1 = 4 (-3,-8) --> x2 = -3, y2 =-8 does that help

OpenStudy (anonymous):

i have to explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did ..... i dont get that part

OpenStudy (dumbcow):

it doesn't matter what point you use

OpenStudy (dumbcow):

whatever you think looks nicer i guess :)

OpenStudy (anonymous):

y - 4 = 2 ( x - 3) ?